# Further results for classical and universal characters twisted by roots of unity

### v1

@@ Abstract.
We revisit factorizations of classical characters under various specializations, some old and some new. We first show that all characters of classical families of groups twisted by odd powers of an even primitive root of unity factorize into products of characters of smaller groups. Motivated by conjectures of Wagh and Prasad (Manuscr. Math. 2020), we then observe that certain specializations of Schur polynomials factor into products of two characters of other groups. We next show, via a detour through hook Schur polynomials, that certain Schur polynomials indexed by staircase shapes factorize into linear pieces. Lastly, we consider classical and universal characters specialized at roots of unity. One of our results, in parallel with Schur polynomials, is that universal characters take values only in $\{0,\pm 1,\pm 2\}$ at roots of unity.
@@ 1. Introduction
The irreducible characters of the classical families of groups possess many remarkable properties. In this work, we focus on the factorization properties of these characters under different specializations building on our previous work [AK22, Kum24, Kum22]. This subject has a venerable history and we refer to an upcoming history by Albion [Alb24] for a detailed account.
Littlewood and Richardson were certainly one of the first who realized that the characters of the general linear group $\operatorname{GL}_{n}(\mathbb{C})$, also known as Schur polynomials and denoted $s_{\lambda}(x_{1},\dots,x_{n})$, specialize to $-1,0,1$ when the variables are taken to be roots of unity [LR34, Theorem IX]; see Theorem 2.2. In the same work, they also generalized this to a version with many variables [LR34, Theorem VI]; see Theorem 2.3. This was later published in Littlewood’s book [Lit06, Chapter 7]. This latter result has been independently rediscovered by Prasad [Pra16], and another proof has recently been given by Karmakar [Kar22]. Kumar [Kum23] has generalized these results to flagged skew Schur polynomials.
In an almost unnoticed work, Lecouvey [Lec09a] generalized these classical character factorizations, which were found independently by the authors [AK22]. Kumari has generalized these factorizations with more specializations [Kum24] and to hook Schur functions [Kum22]. Lecouvey [Lec09b], in another little-known work, also generalized these results to universal characters of other families of Weyl groups of types B, C and D, and these were again found independently by Albion [Alb23]. Very recently, Karmakar [Kar24] has considered the classical characters evaluated at elements of order 2, generalizing a special case of these factorization results.
@@
In this work, we extend several of these factorization results. We first consider the factorizations of classical characters twisted by odd powers of an even primitive root of unity in Section 3. In Section 4, we prove relations among characters of different groups when specialized to roots of unity. In Section 5, we begin by discussing when a universal character of classical type is independent of variables which are twisted by roots of unity. As a consequence, we provide necessary and sufficient conditions for a Schur polynomial in two finite sets of variables to be equal to the corresponding hook Schur polynomial. Lastly, this will prove factorization results for Schur polynomials indexed by staircase shapes. Finally, in Section 6 and Section 7, we will give explicit formulas for the universal and classical characters evaluated at roots of unity respectively.
@@ 2.1. Partitions, cores and quotients
An (integer) partition $\lambda=(\lambda_{1},\dots,\lambda_{n})$ is a weakly decreasing sequence of nonnegative integers. The elements of the partition are called parts. Let $\mathcal{P}$ be the set of all partitions. The size $|\lambda|$ of a partition $\lambda$ is the sum of its parts, and the length $\ell(\lambda)$ of $\lambda$ is the number of nonzero parts. Let $\mathcal{P}_{n}$ be the set of all partitions with length at most $n$.
@@
It is well-known that irreducible (polynomial) representations of $\operatorname{GL}_{n}$ are indexed by partitions in $\mathcal{P}_{n}$. Suppose $\lambda\in\mathcal{P}_{n}$, with $\lambda_{1}=c$ indexes an irreducible representation of $\operatorname{GL}_{n}$. Then the dual representation is indexed by the partition $\hat{\lambda}=(c-\lambda_{n},\dots,c-\lambda_{1})$. In other words, $\hat{\lambda}$ is obtained by considering the complement of the Young diagram of $\lambda$ inside the $c\times n$ rectangle, and reading it in reverse order. For example, if $\lambda=(5,2,1,1)\in\mathcal{P}_{4}$, then $c=5$ and so $\hat{\lambda}=(4,4,3,0)$. If $\lambda$ is such that $\hat{\lambda}=\lambda$, then $\lambda$ is said to index a self-dual representation. Notice that duality is tied to size of the underlying group. For example, $(4,2)$ is self-dual for $\operatorname{GL}_{3}$, but not for $\operatorname{GL}_{4}$.
@@
For a partition $\lambda$ and an integer $m$ such that $\lambda\in\mathcal{P}_{m}$, define the beta-set of $\lambda$ by $\beta(\lambda,m)=(\beta_{1}(\lambda,m),\dots,\beta_{m}(\lambda,m))$ where $\beta_{i}(\lambda,m)=\lambda_{i}+m-i$. For $10\leq i\leq t-$ 1, suppose $n_{i}(\lambda,m)$ is the number of parts of $\beta(\lambda,m)$ congruent to $i\pmod{t}$. We will write $\beta(\lambda)$ and $n_{i}(\lambda)$ whenever $m$ is clear from the context.
For a partition $\lambda\in\mathcal{P}_{tn}$, let $\sigma_{\lambda}\in S_{tn}$ be the permutation that rearranges the parts of $\beta(\lambda)\equiv\beta(\lambda,tn)$ such that
@@
Let $t>2$ be a fixed positive integer throughout the article. Macdonald defines the $t$-core and $t$-quotient of $\lambda$ using its beta-set. Suppose $\beta_{j}^{(i)}(\lambda)$, $1\leq j\leq n_{i}(\lambda)$ be the $n_{i}(\lambda)$ parts of $\beta(\lambda)$ congruent to $i\pmod{t}$ in decreasing order. Arrange the $tn$ numbers $tj+i$, where $10\leq j\leq n_{i}(\lambda)-$ 1 and $10\leq i\leq t-$ 1 in descending order, say $\tilde{\beta}_{1}>\dots>\tilde{\beta}_{tn}$. Then the $t$-core of $\lambda$ has parts $\tilde{\beta}_{i}-tn+i$, $1\leq i\leq tn$, and is denoted by $\mathrm{core}_{t}{(\lambda)}$. Partitions $\lambda$ that satisfy $\mathrm{core}_{t}{(\lambda)}=\lambda$ are called $t$-cores. Let $\mathcal{C}_{t}$ denote the set of all $t$-cores.
@@
For a partition $\lambda$, let $\operatorname{rev}(\lambda)$ denote the tuple formed by reversing the parts of $\lambda$ and $-\lambda$ denote the tuple whose entries are negatives of those of $\lambda$. Further, for $a\in\mathbb{N}$, let $a+\lambda$ denote the partition formed by adding $a$ to every part of $\lambda$. Given partitions $\lambda=(\lambda_{1},\dots,\lambda_{m})$ and $\mu=(\mu_{1},\dots,\mu_{n})$ such that $\lambda_{m}\geq\mu_{1}$, we denote by $(\lambda,\mu)$ the partition obtained by concatenating the parts of $\lambda$ and $\mu$ in that order. Building on this notation, we will often need, for $\lambda,\mu\in\mathcal{P}$ the partition $\mu_{1}+(\lambda,0,\dots,0,-\operatorname{rev}(\mu))$, where the number of 00’s is arbitrary. We will use the notation $(\lambda,-\mu)_{n}$ to denote such a partition, where the total number of 00’s is $n-\ell(\lambda)-\ell(\mu)$. Note that $n$ must be at least as large as $\ell(\lambda)+\ell(\mu)$. For example
@@
Notice that the last element of $(\lambda,-\mu)_{n}$ is always 00. We will also use the notation (±$(\pm\lambda)_{n}=(\lambda,-\lambda)_{n}$ often, which is self-dual for all $n\geq 2\ell(\lambda)$.
@@ 2.2. Classical characters
The classical familes of simple Lie groups are classified into the general linear groups $\operatorname{GL}_{n}$, the orthogonal groups $\operatorname{SO}_{n}$ and the symplectic groups $\operatorname{Sp}_{2n}$. The general linear groups are said to belong to type A, the orthogonal groups of odd size, to type B, the symplectic groups, to type C, and the orthogonal groups of even size, to type D. In each of these cases, the irreducible characters are symmetric polynomials or Laurent polynomials in certain variables, and are indexed by partitions.
@@
The ordinarly generating function of the left hand side is
@@
For a partition $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathcal{P}_{n}$, the Schur polynomial or $\operatorname{GL}_{n}$ character is given by
@@ 2.3. Universal characters
The Schur polynomial is itself a universal character because the $\operatorname{GL}_{n}$ character $s_{\lambda}(X)$, when $\ell(\lambda)\leq n-1$ and $x_{n}$ is set to 00, becomes the $\operatorname{GL}_{n-1}$ character indexed by the same partition $\lambda$. This is not true for the other characters because they are Laurent polynomials in the $x_{i}$’s and so we cannot set $x_{n}$ to 00 there.
Koike and Terada [KT90] defined universal characters to address this issue for classical groups of other types. These characters are symmetric polynomials (not Laurent polnomials) which reduce correctly.
Suppose $X_{\infty}=(x_{1},x_{2},\dots)$. Koike and Terada [KT90] defined the universal character for $\mathrm{O}_{n}$ as
(2.11) $\mathfrak{o}_{\lambda}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})-h_{\lambda_{i}-i-j}(X_{\infty})\right)_{1\leq i\leq n},$
and the universal character for $\operatorname{Sp}_{2n}$,
(2.12) $\mathfrak{sp}_{\lambda}(X_{\infty})=\frac{1}{2}\det\left(h_{\lambda_{i}-i+j}(X_{\infty})+h_{\lambda_{i}-i-j+2}(X_{\infty})\right)_{1\leq i\leq n}.$
@@
Koike [Koi97, Definition 6.4] later defined the universal character for the group $\operatorname{SO}_{2n+1}$. If $\lambda\in\mathcal{P}_{n}$, he set
(2.14) $\mathfrak{so}_{\lambda}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})+h_{\lambda_{i}-i-j+1}(X_{\infty})\right)_{1\leq i\leq n},$
and showed that $\operatorname{so}_{\lambda}(X)=\mathfrak{so}_{\lambda}(X,\overline{X},0,0,\dots)$. Albion [Alb23, Section 2.3] defines a variant of the above as
$\mathfrak{so}_{\lambda}^{-}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})-h_{\lambda_{i}-i-j+1}(X_{\infty})\right)_{1\leq i\leq n},$
and proves that
(2.15) $\mathfrak{so}_{\lambda}^{-}(X_{\infty})=(-1)^{|\lambda|}\mathfrak{so}_{\lambda}(X_{\infty}).$
For partitions $\lambda$ and $\mu$, Koike [Koi89] defined a universal character of $\operatorname{GL}_{n}(\mathbb{C})$ by
@@
In what follows, it should be understood that a universal character with finitely many variables means that all the other variables are set to 00.
@@ 2.4. Useful factorizations
Here we write down all the factorization results we use later in the article. We first begin with those for $\operatorname{GL}_{n}$ characters, i.e. Schur polynomials. Recall that $t\geq 2$ is a positive integer and $\omega$ is a primitive $t$’th root of unity. The evaluation of the Schur polynomials at the roots of unity is astonishingly simple.
@@
Let $\lambda\in\mathcal{P}_{tn}$ be a partition indexing an irreducible representation of $\operatorname{GL}_{tn}$ and $\mathrm{quo}_{t}{(\lambda)}=(\lambda^{(0)},\dots,\lambda^{(t-1)})$. Then the $\operatorname{GL}_{tn}$-character $s_{\lambda}(X,\omega X,\dots,\omega^{t-1}X)$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}$ is empty. In that case, it is given by
@@
Fix $10\leq m\leq t-$ 1. Let $\lambda$ be a partition in $\mathcal{P}_{tn+m}$. Then the $\operatorname{GL}_{tn+m}$-character $s_{\lambda}(X,\omega X,\dots,\omega^{t-1}X,y,\omega y,\dots,\omega^{m-1}y)$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}=\nu\coloneqq(\nu_{1},\dots,\nu_{m})$ with $\nu_{1}\leq t-m$. In that case,
@@
(2.22) $\begin{split}\mathfrak{o}_{\lambda}(X_{\infty},\omega X_{\infty},\dots,\omega^{t-1}X_{\infty})=(-1)^{\epsilon}\operatorname{sgn}(\sigma_{\lambda})\,&\mathfrak{o}_{\lambda^{(0)}}(X_{\infty}^{t})\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}\operatorname{rs}_{\lambda^{(i)},\lambda^{(t-i)}}(X_{\infty}^{t})\\ &\times\begin{cases}\mathfrak{so}^{-}_{\lambda^{(t/2)}}(X_{\infty}^{t})&t\text{ even},\\ 1&t\text{ odd},\end{cases}\end{split}$
@@
(2.23) $\begin{split}\mathrm{o}^{\text{even}}_{\lambda}(X,\omega X,\dots,\omega^{t-1}X,1)=(-1)^{\epsilon}\,\operatorname{sgn}(\sigma_{\lambda})&\,\mathrm{o}^{\text{even}}_{\lambda^{(0)}}(X^{t},1)\prod_{i=1}^{\left\lfloor{(t-1)}/{2}\right\rfloor}s_{(\lambda^{(i)},-\lambda^{(t-i)})_{2n}}(X^{t},\overline{X}^{t})\\ \times&\begin{cases}(-1)^{\sum_{i}\lambda_{i}^{(t/2)}}\operatorname{so}_{\lambda^{(t/2)}}(-X^{t})&t\text{ even,}\\ 1&t\text{ odd,}\end{cases}\end{split}$
@@
We note that specializing the variables to $(X,\overline{X})$ in Theorem 2.5, Theorem 2.6 and Theorem 2.8 gives corresponding factorization results for classical characters [AK22]. The next two results are factorizations for characters of self-dual representations of $\operatorname{GL}_{n}$ when half the variables are reciprocals of the other half.

### v2

@@ Abstract.
We revisit factorizations of classical characters under various specializations, some old and some new. We first show that all characters of classical families of groups twisted by odd powers of an even primitive root of unity factorize into products of characters of smaller groups. Motivated by conjectures of Prasad and Wagh (Manuscr. Math. 2022), we then observe that certain specializations of Schur polynomials factor into products of two characters of other groups. We next show, via a detour through hook Schur polynomials, that certain Schur polynomials indexed by staircase shapes factorize into linear pieces. Lastly, we consider classical and universal characters specialized at roots of unity. One of our results, in parallel with Schur polynomials, is that universal characters take values only in $\{0,\pm 1,\pm 2\}$ at roots of unity.
@@ 1. Introduction
The irreducible characters of the classical families of groups possess many remarkable properties. In this work, we focus on the factorization properties of these characters under different specializations building on our previous work [AK22, Kum24]. This subject has a venerable history and we refer to Albion [Alb25] for a detailed account.
Littlewood and Richardson were certainly one of the first who realized that the characters of the general linear group $\operatorname{GL}(n)$, also known as Schur polynomials and denoted $s_{\lambda}(x_{1},\dots,x_{n})$, specialize to 1,0-1,0 or $1$ when the variables are taken to be roots of unity [LR34, Theorem IX]; see Theorem 2.2. In the same work, they also generalized this to a version with many variables [LR34, Theorem VI]; see Theorem 2.3. This was later published in Littlewood’s book [Lit06, Chapter 7]. This latter result has been independently rediscovered by Prasad [Pra16], and another proof has recently been given by Karmakar [Kar22]. Kumar [Kum23] has generalized these results to flagged skew Schur polynomials.
In an almost unnoticed work, Lecouvey [Lec09a] generalized these classical character factorizations, which were found independently by the authors [AK22]. Kumari has generalized these factorizations with more specializations [Kum24]. Lecouvey [Lec09b], in another little-known work, also generalized these results to universal characters of other families of Weyl groups of types B, C and D, and these were again found independently by Albion [Alb23]. Very recently, Karmakar [Kar24] has considered the classical characters evaluated at elements of order 2, generalizing a special case of these factorization results, and Nadimpalli–Pattanayak–Prasad [NPP25] have given representation-theoretic proofs of some of these results.
@@
In this work, we extend several of these factorization results. Our motivation is to find new factorizations and document some of those not yet written down in the literature. We first consider the factorizations of classical characters twisted by odd powers of an even primitive root of unity in Section 3. That this special case would have nice factorizations was not clear beforehand, and this may be a case that may be amenable to representation-theoretic tools. In Section 4, we prove unexpected relations among characters of different groups when specialized to roots of unity. Although most of the factorizations here are not new, what is new is that they are related to each other in some very simple way. In Section 5, we begin by discussing when a universal character of classical type is independent of variables which are twisted by roots of unity. Here, we note a phenomenon we have not seen before, namely that certain symmetric polynomials become independent of some of the variables. As a consequence, we provide necessary and sufficient conditions for a Schur polynomial in two finite sets of variables to be equal to the corresponding hook Schur polynomial. This will also prove factorization results for Schur polynomials indexed by staircase shapes. Finally, in Section 6 and Section 7, we will give explicit formulas for the universal and classical characters evaluated at roots of unity respectively. While it is true that these latter results are special cases of earlier older results, it is helpful to write them down separately. For instance, it is not apriori clear that $\mathfrak{o}_{\lambda}$ evaluated at roots of unity is either 0 or $\pm 1$ (Theorem 6.2) .
@@ 2.1. Partitions, cores and quotients
An (integer) partition $\lambda=(\lambda_{1},\dots,\lambda_{n})$ is a weakly decreasing sequence of nonnegative integers. The elements of the partition are called parts. Let $\mathcal{P}$ be the set of all partitions. The size $|\lambda|$ of a partition $\lambda$ is the sum of its parts, and the length $\ell(\lambda)$ of $\lambda$ is the number of nonzero parts. Let $\mathcal{P}_{n}$ be the set of all partitions with length at most $n$. For $\lambda\in\mathcal{P}_{n}$ and $a\geq 0$, we denote the partition $(a+\lambda_{1},\dots,a+\lambda_{n})$ by $a+\lambda$.
@@
It is well-known that irreducible (polynomial) representations of $\operatorname{GL}(n)$ are indexed by partitions in $\mathcal{P}_{n}$. Suppose $\lambda\in\mathcal{P}_{n}$, with $\lambda_{1}=c$ indexes an irreducible representation of $\operatorname{GL}(n)$. Then the dual representation is indexed by the partition $\hat{\lambda}=(c-\lambda_{n},\dots,c-\lambda_{1})$. In other words, $\hat{\lambda}$ is obtained by considering the complement of the Young diagram of $\lambda$ inside the $c\times n$ rectangle, and reading it in reverse order. For example, if $\lambda=(5,2,1,1)\in\mathcal{P}_{4}$, then $c=5$ and so $\hat{\lambda}=(4,4,3,0)$. If $\lambda$ is such that $\hat{\lambda}=\lambda$, then $\lambda$ is said to index a self-dual representation. Notice that duality is tied to the rank of the underlying group. For example, $(4,2)$ is self-dual for $\operatorname{GL}(3)$, but not for $\operatorname{GL}(4)$.
@@
For a partition $\lambda$ and an integer $m$ such that $\lambda\in\mathcal{P}_{m}$, define the beta set of $\lambda$ by $\beta(\lambda,m)=(\beta_{1}(\lambda,m),\dots,\beta_{m}(\lambda,m))$ where $\beta_{i}(\lambda,m)=\lambda_{i}+m-i$. Let $t>2$ be a fixed positive integer throughout the article. For $0\leq i\leq t-1$, suppose $n_{i}(\lambda,m)$ is the number of parts of $\beta(\lambda,m)$ congruent to $i\pmod{t}$. We will write $\beta(\lambda)$ and $n_{i}(\lambda)$ whenever $m$ is clear from the context.
For a partition $\lambda\in\mathcal{P}_{tn}$, let $\sigma_{\lambda}\in S_{tn}$ be the permutation that rearranges the parts of $\beta(\lambda)=\beta(\lambda,tn)$ such that
@@
Macdonald defines the $t$-core and $t$-quotient of $\lambda$ using its beta set. Suppose $\beta_{j}^{(i)}(\lambda)$, $1\leq j\leq n_{i}(\lambda)$ are the $n_{i}(\lambda)$ parts of $\beta(\lambda)$ congruent to $i\pmod{t}$ in decreasing order. Arrange the $tn$ numbers $tj+i$, where $0\leq j\leq n_{i}(\lambda)-1$ and $0\leq i\leq t-1$ in descending order, say $\tilde{\beta}_{1}>\dots>\tilde{\beta}_{tn}$. Then the $t$-core of $\lambda$ has parts $\tilde{\beta}_{i}-tn+i$, $1\leq i\leq tn$, and is denoted by $\mathrm{core}_{t}{(\lambda)}$. Partitions $\lambda$ that satisfy $\mathrm{core}_{t}{(\lambda)}=\lambda$ are called $t$-cores. Let $\mathcal{C}_{t}$ denote the set of all $t$-cores.
@@
For a partition $\lambda$, let $\operatorname{rev}(\lambda)$ denote the tuple formed by reversing the parts of $\lambda$ and $\lambda$ denote the tuple whose entries are negatives of those of $\lambda$. Further, for $a\in\mathbb{N}$, let $a+\lambda$ denote the partition formed by adding $a$ to every part of $\lambda$. Given partitions $\lambda=(\lambda_{1},\dots,\lambda_{m})$ and $\mu=(\mu_{1},\dots,\mu_{n})$ such that $\lambda_{m}\geq\mu_{1}$, we denote by $(\lambda,\mu)$ the partition obtained by concatenating the parts of $\lambda$ and $\mu$ in that order. Building on this notation, we will often need, for $\lambda,\mu\in\mathcal{P}$ the partition $\mu_{1}+(\lambda,0,\dots,0,-\operatorname{rev}(\mu))$, where the number of 0’s is arbitrary. We will use the notation $(\lambda,-\mu)_{n}$ to denote such a partition, where the total number of 0’s is $n-\ell(\lambda)-\ell(\mu)$. Note that $n$ must be at least as large as $\ell(\lambda)+\ell(\mu)$. For example
@@
Notice that the last element of $(\lambda,-\mu)_{n}$ is always 0. We will also use the notation $(\pm\lambda)_{n}=(\lambda,-\lambda)_{n}$ often, which is self-dual for all $n\geq 2\ell(\lambda)$.
@@ 2.2. Classical characters
The classical families of simple Lie groups are classified into the general linear groups $\operatorname{GL}(n)$, the orthogonal groups $\operatorname{SO}(n)$ and the symplectic groups $\operatorname{Sp}(2n)$. The general linear groups are said to belong to type A, the orthogonal groups of odd size, to type B, the symplectic groups, to type C, and the orthogonal groups of even size, to type D. In each of these cases, the irreducible characters are symmetric polynomials or Laurent polynomials in certain variables, and are indexed by partitions.
@@
The ordinary generating function of the left hand side is
@@
For a partition $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathcal{P}_{n}$, the Schur polynomial or $\operatorname{GL}(n)$ character is given by
@@ 2.3. Universal characters
The Schur polynomial is itself a universal character because the $\operatorname{GL}(n)$ character $s_{\lambda}(X)$, when $\ell(\lambda)\leq n-1$ and $x_{n}$ is set to 0, becomes the $\operatorname{GL}({n-1})$ character indexed by the same partition $\lambda$. This is not true for the other characters because they are Laurent polynomials in the $x_{i}$’s and so we cannot set $x_{n}$ to 0 there.
Koike and Terada [KT90] defined universal characters (see below) to address this issue for classical groups of other types. These characters are symmetric polynomials (not Laurent polynomials) which reduce correctly.
Suppose $X_{\infty}=(x_{1},x_{2},\dots)$. Koike and Terada [KT90] defined the universal character for $\mathrm{O}({n})$ as
(2.11) $\mathfrak{o}_{\lambda}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})-h_{\lambda_{i}-i-j}(X_{\infty})\right)_{{1\leq i,j\leq n}},$
and the universal character for $\operatorname{Sp}({2n})$,
(2.12) $\mathfrak{sp}_{\lambda}(X_{\infty})=\frac{1}{2}\det\left(h_{\lambda_{i}-i+j}(X_{\infty})+h_{\lambda_{i}-i-j+2}(X_{\infty})\right)_{{1\leq i,j\leq n}}.$
@@
Koike [Koi97, Definition 6.4] later defined the universal character for the group $\operatorname{SO}({2n+1})$. If $\lambda\in\mathcal{P}_{n}$, he set
(2.14) $\mathfrak{so}_{\lambda}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})+h_{\lambda_{i}-i-j+1}(X_{\infty})\right)_{{1\leq i,j\leq n}},$
and showed that $\operatorname{so}_{\lambda}(X)=\mathfrak{so}_{\lambda}(X,\overline{X},0,0,\dots)$. Albion [Alb23, Section 2.3] defined a variant of the above as
(2.15) $\mathfrak{so}_{\lambda}^{-}(X_{\infty})=\det\left(h_{\lambda_{i}-i+j}(X_{\infty})-h_{\lambda_{i}-i-j+1}(X_{\infty})\right)_{{1\leq i,j\leq n}},$
and proved that
(2.16) $\mathfrak{so}_{\lambda}^{-}(X_{\infty})=(-1)^{|\lambda|}\mathfrak{so}_{\lambda}(-X_{\infty}).$}
For partitions $\lambda$ and $\mu$, Koike [Koi89] defined a universal character of $\operatorname{GL}(n)$ by
@@
In what follows, it should be understood that a universal character with finitely many variables means that all the other variables are set to 0.
@@ 2.4. Useful factorizations
Here we write down all the factorization results we use later in the article. We first begin with those for $\operatorname{GL}(n)$ characters, i.e. Schur polynomials. Recall that $t\geq 2$ is a positive integer and $\omega$ is a primitive $t$’th root of unity. The evaluation of the Schur polynomials at the roots of unity is astonishingly simple.
@@
Let $\lambda\in\mathcal{P}_{tn}$ be a partition indexing an irreducible representation of $\operatorname{GL}({tn})$ and $\mathrm{quo}_{t}{(\lambda)}=(\lambda^{(0)},\dots,\lambda^{(t-1)})$. Then the $\operatorname{GL}({tn})$-character $s_{\lambda}(X,\omega X,\dots,\omega^{t-1}X)$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}$ is empty. In that case, it is given by
@@
Fix $0\leq m\leq t-1$. Let $\lambda$ be a partition in $\mathcal{P}_{tn+m}$. Then the $\operatorname{GL}({tn+m})$-character $s_{\lambda}(X,\omega X,\dots,\omega^{t-1}X,y,\omega y,\dots,\omega^{m-1}y)$ is nonzero if and only if $\mathrm{core}_{t}{(\lambda)}=\nu\coloneqq(\nu_{1},\dots,\nu_{m})$ with $\nu_{1}\leq t-m$. In that case,
@@
[…]
@@
[…]
@@
We note that specializing the variables to $(X,\overline{X})$ in Theorem 2.5, Theorem 2.6 and Theorem 2.8 gives corresponding factorization results for classical characters [AK22]. The next two results are factorizations for characters of self-dual representations of $\operatorname{GL}(n)$ when half the variables are reciprocals of the other half.